# These $\equalcontThese authors contributed equally to this work.$

### v1

@@ 3.3 Re-embedded features aggregation
Re-embedded features aggregation is an essential module of our foundation model fusion framework. The multi-instance learning (MIL) is widely adopted as an effective solution in WSI analysis, where the labels $\mathbf{y}$ are assigned with each WSI. And the WSI $X$ can be defined as bag, the patches within $X$ are instances. To effectively demonstrate the features aggregation, we will first briefly introduce MIL and then proceed to the features aggregation in our FuseCPath.
The multi-instance learning (MIL) is a useful weakly supervised learning method in WSI analysis. In MIL, the training set consists of several bags with labels $\mathbf{y}=\{y_{1},y_{2},...,y_{D}\}$, and each bag contains several instances without labels. If at least one instance in a bag is positive, the bag is considered as a positive bag; if all instances in a bag are negative instances, the bag is considered a negative bag. Considering the situation in binary classification, we define $B=\left\{(x_{1},y_{1}),...,(x_{D},y_{D})\right\}$ as a bag. $x_{d}\ (d\in\{1,2,...,D\})$ are instances with labels $y_{d}\in{0,1}$, the label $Y$ of $B$ is given by:
@@
In this work, the features aggregation $\mathcal{A}(\cdot)$ is implemented by Attention-based multi-instances learning (AB-MIL), which integrates the strengths of attention-based MIL pooling for aggregating the features $\hat{\mathbf{Z}}^{pe}$ into a single feature vector $\mathbf{Z}^{pe}$ with a weighted averaging operation. Figure 1 (c) demonstrates the procedure of features aggregation via AB-MIL and $\mathbf{Z}^{pe}$. In this work, the input of AB-MIL is the re-embedded patch-level FMs features $\hat{\mathbf{Z}}^{pe}\in\mathbb{R}^{(N_{K}\times K)\times D}$. By adopting AB-MIL as pooling module, the aggregated feature $\mathbf{Z}^{pe}$ can be formulated as follows:
@@
$a_{d}(\hat{\mathbf{Z}}^{pe}_{d})=\frac{\exp\left({\mathbf{W}^{T}\tanh\left(\mathbf{V}(\hat{\mathbf{Z}}^{pe}_{d})^{T}\right)}\right)}{\sum_{j=1}^{D}\exp\left({\mathbf{W}^{T}\tanh\left(\mathbf{V}(\hat{\mathbf{Z}}^{pe}_{j})^{T}\right)}\right)},$
@@ 3.4 Slide-level collaborative distillation
Slide-level foundation model is capable of yielding a high-level representation of WSI, which is more concise for a downstream task fine-tuning. However, it is a problem that fuse these slide-level global representations with patch-level local representations. This is challenging due to the dimensional gaps. In the proposed FuseCPath, we try to solve this problem by regarding the slide-level features as soft labels derived from teacher models. Consequently, we propose a slide-level collaborative distillation strategy to fuse slide-level FMs that contain global features simultaneously.
@@
$\mathbf{h}^{i}_{se}=\text{Linear}_{i}(\mathbf{L}^{i}_{se}),\mathbf{L}^{i}_{se}\in\mathbb{R}^{1\times d^{i}_{se}}$ (10)
@@
where $C$ denotes the category length of classification-related tasks, such as cancer sub-typing, grading, and biomarker prediction. Given label $\mathbf{y}$, the combined loss function of distillation is formulated as follows:
$\mathcal{L}_{fuse}=\lambda\mathcal{L}_{task}(\mathbf{h}_{pe},\mathbf{y})+(1-\lambda)\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}^{i}_{dist},$ (14)
@@ 7 $\mathbf{Z}^{pe}\leftarrow\text{ReEmbeddedFeaturesAggregation}(\hat{\mathbf{Z}}^{pe})$;
9 $\hat{\mathbf{y}}\leftarrow\text{TaskHeader}(\mathbf{Z}^{f})$;
@@ 4.1 Dataset description and evaluation metrics
$\text{MSE}(\mathbf{y}_{pred},\mathbf{y})=\frac{1}{N}\sqrt{\sum^{N=10}_{i=1}\|\mathbf{y}^{i}_{pred}-\mathbf{y}^{i}\|^{2}}.$ (15)
@@
$C_{\text{index}}=\frac{1}{n}\sum_{i\in\{i,...,n|\delta_{i}=1\}}\sum_{t_{i}>t_{j}}I\left[f_{i}>f_{j}\right],$ (16)
@@
Metrics for clustering. Due to the ground truth labels for clustering evaluation are unavailable, we select two widely adopted metrics for evaluating the clustering, which are the silhouette coefficient (SC) and the Calinski-Harabasz (CH) index. The value of SC ranges from -1 to 1, where values approaching 0 suggest cluster overlap, negative values indicate incorrect sample assignments, and higher positive values reflect well-separated clusters. The CH index evaluates clustering quality by calculating the ratio of between-cluster variance to within-cluster variance, where variance is defined as the sum of squared Euclidean distances. Higher CH values indicate better clustering results, reflecting both strong separation between different clusters and high compactness within individual clusters.
@@ 4.2 Biomarker predictions
To quantitatively evaluate the performance of each method, we present the results assessed by AUROC in Table 2. The experimental results demonstrate that the proposed FuseCPath framework outperforms these baseline MIL-based methods. Performance improvements are observed across multiple biomarker prediction tasks on datasets TCGA-LUAD and TCGA-BLCA, with an average increase of 12.7% in AUROC compared to the baseline methods with the best performance. The superior results can be attributed to the teacher model’s high-level feature representations, which provide additional discriminative information to guide the student models’ feature fusion process. These results prove that our re-embedding and distillation-based FuseCPath framework enhances feature learning, particularly in scenarios with class imbalance and limited labeled training data.
@@
The comprehensive evaluation results are presented in Table 3, which is measured by AUROC, reveal several key findings: First, FuseCPath consistently outperforms all individual FMs across all three prediction tasks. The observed average performance has improved by 17%. This substantial improvement can be attributed to two fundamental advantages of our FuseCPath framework: First, the effective fusion of complementary features from heterogeneous FMs through our proposed re-embedding and distillation mechanism. Second, the patch-level and slide-level collaborative fusion of FMs adaptively emphasizes the most meaningful features for the prediction of each biomarker. These results imply that an effective fusion of diverse FMs can yield superior predictive capability compared to a single model, as the ensemble approach mitigates individual model limitations while preserving their respective strengths through feature re-embedding and distillation. The results also prove the importance of the fusion of pathology FMs, demonstrating that a carefully devised fusion framework can improve the model’s performance by leveraging the rich but complementary information contained in diverse FMs.
@@ 4.3 Gene expression prediction
In this experiment, we evaluate the performance of gene expression prediction for our method. The proposed FuseCPath is capable of predicting the expression of many genes involved in pathways. We select 10 popularly investigated genes to assess the prediction errors of the proposed FuseCPath, which are TP53, EGFR, KRAS, BRAF, PIK3CA, IDH1, FGFR3, RB1, ATM, and ERBB2. Expression are evaluated by logarithmically transformed transcripts per million (TPM) values $t$, which are formulated as follows:
@@
We select two methods for our comparisons, which are the proposed FuseCPath (wdist) and the method only containing patch-level features using R2T without the distillation module (w/o dist). Figure 5 presents radar charts comparing predicted and ground truth gene expression, visually illustrating the alignment between our model’s predictions and ground truths. We present the quantitative results averaged across all samples from the validation datasets of TCGA-LUAD, TCGA-BLCA, and TCGA-COAD in Table 4, respectively. To evaluate the prediction error, we provide the mean square error (MSE) comparisons between the prediction results of different methods and the ground truth in Table 5, demonstrating the accuracy of the expression for individual genes.
@@
4.4 Survival Analysis
In this experiment, we evaluate and compare the performance of the proposed FuseCPath with the features from several slide-level FMs, which are CHIEF [12], Gigapath-SE [14], and PRISM [11]. We provide the results of Kaplan-Meier survival curves for each comparison method in Figure 6. The test cohorts are divided into high- and low-risk groups using the median risk score predicted by our proposed FuseCPath framework. Comparative analysis demonstrates that FuseCPath achieves significantly improved risk stratification, producing a more distinct separation between the two risk groups with enhanced prognostic discrimination capability. This implies that the FuseCPath consistently performs better the the other FMs in distinguishing high- and low-risk patients.
@@ 4.5 Analysis of clustering
Before the training process of the FuseCPath, one essential step is to select representative image patches from the input WSIs. In this work, to integrate the features from heterogeneous patch-level FMs, we devise a multi-view clustering-based strategy to partition the patches into $K=50$ clusters. Each embedding from the corresponding foundation model can be regarded as a view of the WSI. In this section, we conduct an experimental analysis of patch clustering.
@@
Visualization and interpretability. We present the visualized results of the patch clustering for each example WSI in Figure 7. To better demonstrate the visualization results, we present the zoom-in areas in original WSIs alongside their corresponding clustering results. The results clearly show that under the guidance of embeddings from the heterogeneous FMs, the clustering results exhibit clear alignment with cellular morphological distributions. The clustered regions are related to tissue structures, indicating that the multi-view clustering can capture both local and global morphological patterns. This implies that the selected patches are representative and reliable enough for the training of FuseCPath.
Analysis of the number of clusters. To determine the optimal number of clusters ($K$) for our multi-view clustering, we performed a systematic evaluation guided by both quantitative metrics and biological considerations. We ultimately selected $K=50$ as it demonstrated the best balance between cluster separation and cohesion. It was informed by the known biological diversity of human cells, with a single-cell atlas identifying 102 distinct cell types, leading us to set an upper bound of 110 clusters to maintain biological relevance. To ensure an adequate representation of histological patterns, we set a lower bound of 30 clusters in this experiment. As fewer clusters probably affect the diversity of selected patches. We validate this range by evaluating the quality of the cluster at 10-cluster intervals in Table 7. The selected metrics are the silhouette coefficient (SC) and Calinski-Harabasz (CH). From the results, we can conclude that $K=50$ emerges as the optimal choice that satisfies both our computational metrics and the visual constraints of biological tissues.
@@
Spectral Clustering 0.07 205.3
Agglomerative Clustering 0.07 892.5
@@
Multi-view clustering vs. single-view clustering. In this part, we devise experiments to compare multi-view clustering with other single-view clustering. We select spectral clustering, agglomerative clustering, and affinity propagation for comparisons. These selected methods are clustered with embeddings from CONCH [9]. We present SC and CH quantitative results in Table 8. The results prove that multi-view clustering achieves a higher quality of clustering compared to the other single-view clustering methods. Multi-view clustering integrates complementary features from heterogeneous patch-level FMs into a unified representation space, whereas single-view clustering, such as spectral clustering, operates on a single feature space. ConsequentlyAdditionally, mMulti-view clustering will capture higher-level relationships between different feature spaces, enabling nonlinear pattern discovery beyond single-view clustering limitations.
@@ 4.6 Ablation studies
To evaluate the effectiveness of the main components of our proposed FuseCPath framework, we conduct ablation studies on the following aspects: the number of slide-level FMs and the number of selected patches. All experiments were conducted ton the prediction task of the biomarker TP53 for the TCGA-LUAD and TCGA-BLCA datasets.
@@
Ablation studies on the number of slide-level FMs. In this experiment, we conduct an ablation study on the number of slide-level FMs for distillation. We present the quantitative results of AUROC in Table 9 and Figure 8. The best prediction performance (AUROC) is obtained when 3 slide-level FMs are utilized. The performance is decreasing when the number of slide-level FMs decreases. When FuseCPath is only trained with re-embedded patch-level features, and the performance decreases by 6.5% on average. These results imply that more slide-level FMs utilized for distillation will provide more useful semantic information during training.
@@ 5 Discussion and limitation
In this work, we have proposed a novel framework called FuseCPath for the fusion of heterogeneous pathology FMs from patch-level and slide-level simultaneously. The proposed FuseCPath framework includes the following essential modules to effectively fuse the pathological foundation models: representative patches selection based on multi-view patch features clustering, patch-level features re-embedding & aggregation, and slide-level collaborative knowledge distillation. These modules contribute to the performance improvement of biomarker prediction, gene expression, and survival analysis on datasets TCGA-LUAD, TCGA-BLCA, and TCGA-COAD. In conclusion, The FuseCPath framework will yield a new ensemble model with superior performance in many meaningful downstream tasks like biomarker predictions, gene expression predictions, and survival analysis. In addition, clustering with multi-view features will provide insight into the visualization analysis of tissue morphography.
Despite the demonstrated utility in this article, our proposed FuseCPath poses potential limitations in its capacity to integrate more high-dimensional multi-omics data, such as the emerging spatial transcriptomic technologies. The current framework may not fully capture the underlying non-linear relationships between different omics data. The rapid evolution of FMs presents a promising strategy for the integration of multi-omics data and WSIs [27]. In future research, we will extend the FuseCPath framework to the fusion of foundation model-agnostic gene representations and embeddings from multi-omics FMs to improve the precision of molecular-level WSI analysis.
@@
Zhidong Yang: Methodology, Investigation, Writing - original draft & editing. Xiuhui Shi: Conceptualization, Resources, Data curation. Wei Ba: Conceptualization, Data curation. Zhigang Song: Conceptualization, Data curation. Haijing Luan: Methodology, Validation. Taiyuan Hu: Methodology, Validation. Senlin Lin: Conceptualization, Validation. Jiguang Wang: Writing – review & editing. Shaohua Kevin Zhou: Resources, Writing - review & editing. Rui Yan: Methodology, Resources, Writing - review & editing.

### v2

@@ 3.3 Re-embedded features aggregation
Re-embedded features aggregation is an essential module of our heterogeneous FMs fusion framework. The multi-instance learning (MIL) is widely adopted as an effective solution in WSI analysis, where the labels $\mathbf{y}$ are assigned to each WSI. And the WSI $X$ can be defined as bag, the patches within $X$ are instances. To effectively demonstrate the features aggregation, we will first briefly introduce MIL and then proceed to the features aggregation in our FuseCPath.
The multi-instance learning (MIL) is a useful weakly supervised learning method in WSI analysis. The formulation of MIL is shown as follows. The dataset consists of bags assigned with labels $\mathbf{y}=\{y_{1},y_{2},...,y_{D}\}$. Each bag contains several instances. If at least one instance in a bag is positive, the bag will be considered positive, and if all instances in a bag are negative, the bag will be considered negative. Take the situation in binary classification as an example, we define $B=\left\{(x_{1},y_{1}),...,(x_{D},y_{D})\right\}$ as a bag. $x_{d}\ (d\in\{1,2,...,D\})$ are instances of $B$, each with labels $y_{d}\in{0,1}$. Consequently, the label $Y$ of $B$ is given by:
@@
In this work, the features aggregation $\mathcal{A}(\cdot)$ is implemented by Attention-based multi-instances learning (AB-MIL), which integrates the strengths of attention-based MIL pooling for aggregating the features $\hat{\mathbf{Z}}^{pe}$ into a single feature vector $\mathbf{Z}^{pe}$ with a weighted averaging operation. Figure 2(c) demonstrates the procedure of feature aggregation via AB-MIL. In this work, the input of AB-MIL is the re-embedded patch-level FMs features $\hat{\mathbf{Z}}^{pe}\in\mathbb{R}^{(N_{K}\times K)\times D}$. By adopting AB-MIL as pooling module, the aggregated feature $\mathbf{Z}^{pe}$ can be formulated as follows:
@@
$a_{d}(\hat{\mathbf{Z}}^{pe}_{d})=\frac{\exp\left({\mathbf{W}^{T}\tanh\left(\mathbf{V}(\hat{\mathbf{Z}}^{pe}_{d})^{T}\right)}\right)}{\sum_{j=1}^{D}\exp\left({\mathbf{W}^{T}\tanh\left(\mathbf{V}(\hat{\mathbf{Z}}^{pe}_{j})^{T}\right)}\right)}.$
@@ 3.4 Slide-level collaborative distillation
Slide-level foundation model is capable of yielding a high-level representation of WSI, which is more concise for a downstream task fine-tuning. However, it is a problem that fuse these slide-level global representations with patch-level local representations. This is challenging due to the dimensional gaps. In the proposed FuseCPath, we try to solve this problem by regarding the slide-level features as soft labels derived from teacher models. Consequently, we propose a slide-level collaborative distillation strategy to fuse slide-level FMs that contain global information simultaneously.
@@
$\mathbf{h}^{i}_{se}=\text{Linear}_{i}(\mathbf{L}^{i}_{se}),\mathbf{L}^{i}_{se}\in\mathbb{R}^{1\times d^{i}_{se}},$ (10)
@@
where $C$ denotes the dimension of the softened distributions. Given label $\mathbf{y}$, the combined loss function of distillation is formulated as follows:
$\mathcal{L}_{fuse}=\lambda\mathcal{L}_{task}(\mathbf{h}_{pe},\mathbf{y})+(1-\lambda)\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}^{i}_{dist}.$ (14)
@@ 7 $\mathbf{Z}^{pe}\leftarrow\text{ReEmbeddedFeaturesAggregation}(\hat{\mathbf{Z}}^{pe})$;
9 $\hat{\mathbf{y}}\leftarrow\text{TaskHeader}(\mathbf{Z}^{f};\mathbf{y})$;
@@ 4.1 Dataset description and evaluation metrics
$\text{MSE}(\mathbf{y}_{pred},\mathbf{y})=\frac{1}{N}\sqrt{\sum^{N=10}_{i=1}\|\mathbf{y}^{i}_{pred}-\mathbf{y}^{i}\|^{2}},$ (15)
@@
$C_{\text{index}}=\frac{1}{n}\sum_{i\in\{i,...,n|\delta_{i}=1\}}\sum_{t_{i}>t_{j}}I\left[f_{i}>f_{j}\right].$ (16)
@@
Metrics for clustering. Due to the ground truth labels for the evaluation of clustering being unavailable, we select two widely adopted metrics for evaluating the clustering, which are the silhouette coefficient (SC) and the Calinski-Harabasz (CH) index. The value of SC ranges from -1 to 1, where values approaching 0 suggest cluster overlap, negative values indicate incorrect assignments of the clusters, and higher positive values reflect well-separated clusters. The CH index evaluates clustering quality by calculating the ratio of between-cluster variance to within-cluster variance, where variance is defined as the sum of squared Euclidean distances. Higher CH values indicate better clustering results, reflecting both strong separation between different clusters and high compactness within individual clusters.
@@ 4.2 Biomarker predictions
To quantitatively evaluate the performance of each method, we present the results assessed by AUROC in Table 2. The experimental results demonstrate that the proposed FuseCPath framework outperforms these baseline MIL-based methods. Performance improvements are observed across multiple biomarker prediction tasks on datasets TCGA-LUAD and TCGA-BLCA, with an average increase of 5% in AUROC compared to the baseline methods with the best performance. The superior results can be attributed to the teacher model’s high-level feature representations, which provide additional discriminative information to guide the student models’ feature fusion process. These results prove that our re-embedding and distillation-based FuseCPath framework enhances feature learning, particularly in scenarios with class imbalance and limited labeled training data.
@@
The comprehensive evaluation results are presented in Table 3, which is measured by AUROC, reveal several key findings: First, FuseCPath consistently outperforms all individual FMs across the prediction tasks on TCGA-COAD. The observed average performance has improved by 17%. This improvement can be attributed to two fundamental advantages of our FuseCPath framework: First, the effective fusion of complementary features from heterogeneous FMs through our proposed re-embedding and distillation mechanism. Second, the patch-level and slide-level simultaneous fusion of multiple FMs adaptively emphasizes the most meaningful features for the prediction of each biomarker. These results imply that an effective fusion of diverse FMs can yield superior predictive capability compared to a single model, as the ensemble approach mitigates individual model limitations while preserving their respective strengths through feature fusion. The results also prove the importance of the fusion of pathology FMs, demonstrating that a carefully devised fusion framework can improve the model’s performance by leveraging the rich but complementary information contained in diverse FMs.
@@ 4.3 Gene expression prediction
In this experiment, we evaluate the performance of gene expression prediction for our method. The proposed FuseCPath is capable of predicting the expression of many genes involved in pathways. We select 10 popularly investigated genes to assess the prediction errors of the proposed FuseCPath, which are TP53, EGFR, KRAS, BRAF, PIK3CA, IDH1, FGFR3, RB1, ATM, and ERBB2. Expressions are evaluated by logarithmically transformed transcripts per million (TPM) values $t$, which are formulated as follows:
@@
We select two methods for our comparisons, which are the proposed FuseCPath (wdist) and the method only containing patch-level features using R2T without the distillation module (w/o dist). Figure 5 presents radar charts comparing predicted and ground truth gene expression, visually illustrating the alignment between our model’s predictions and ground truths. We present the quantitative results averaged across all samples from the validation datasets of TCGA-LUAD, TCGA-BLCA, and TCGA-COAD in Table 4, respectively. To evaluate the prediction error, we provide the mean square error (MSE) comparisons between the prediction results of different methods and the ground truth in Table 5, demonstrating the accuracy of the prediction for individual genes.
@@
4.4 Survival Analysis
In this experiment, we evaluate and compare the performance of the proposed FuseCPath with the features from several slide-level FMs, which are CHIEF [12], Gigapath-SE [14], and PRISM [11]. We provide the results of Kaplan-Meier survival curves for each comparison method in Figure 6. The test cohorts are divided into high- and low-risk groups using the median risk score predicted by our proposed FuseCPath framework. Comparative analysis demonstrates that FuseCPath achieves significantly improved risk stratification, producing a more distinct separation between the two risk groups with enhanced prognostic discrimination capability. This implies that the FuseCPath consistently performs better than the other FMs in distinguishing high- and low-risk patients.
@@ 4.5 Analysis of clustering
Before the training process of the FuseCPath, one essential step is to select representative image patches from the input WSIs. In this work, to integrate the features from heterogeneous patch-level FMs, we devise a multi-view clustering-based strategy to partition the patches into $K=50$ clusters. Each embedding from the corresponding foundation model can be regarded as a view of the WSI. In this section, we conduct an experimental analysis of clustering.
@@
Visualization and interpretability. We present the visualized results of the clustering for each example WSI in Figure 7. To better demonstrate the visualization results, we present the zoom-in areas in original WSIs alongside their corresponding clustering results. The results clearly show that under the guidance of embeddings from the heterogeneous FMs, the clustering results exhibit clear alignment with cellular morphological distributions. The clustered regions are related to tissue structures, indicating that the multi-view clustering can capture both local and global morphological patterns. This implies that the patches selected from clusters are representative and reliable enough for the training of FuseCPath.
@@
Spectral Clustering 0.07 209.1
Agglomerative Clustering 0.07 893.3
@@
Analysis of the number of clusters. To determine the optimal number of clusters ($K$) for our multi-view clustering, we performed a systematic evaluation guided by both quantitative metrics and biological considerations. To ensure an adequate representation of histological patterns, we set a lower bound of 30 clusters in this experiment. As fewer clusters probably affect the diversity of selected patches. We validate this range by evaluating the quality of the cluster at 10-cluster intervals in Table 7. The selected metrics are the silhouette coefficient (SC) and Calinski-Harabasz (CH). From the results, we can conclude that $K=50$ emerges as the optimal choice that satisfies both our computational metrics and the visual constraints of biological tissues. So we ultimately selected $K=50$ clusters for the training of FuseCPath.
Multi-view clustering vs. single-view clustering. In this part, we devise experiments to compare multi-view clustering with other single-view clustering. We select spectral clustering, agglomerative clustering, and affinity propagation for comparisons. These selected methods are clustered with embeddings from CONCH [9]. We present SC and CH quantitative results in Table 8. The results prove that multi-view clustering achieves a higher quality of clustering compared to the other single-view clustering methods. Multi-view clustering integrates complementary features from heterogeneous patch-level FMs into a unified representation space, whereas single-view clustering, such as spectral clustering, only operates on a single feature space. Consequently, multi-view clustering will capture higher-level relationships between different feature spaces, enabling nonlinear pattern discovery beyond single-view clustering limitations.
@@ 4.6 Ablation studies
To evaluate the effectiveness of the main components of our proposed FuseCPath framework, we conduct ablation studies on the following aspects: the number of slide-level FMs and the number of selected patches. All experiments were conducted on the prediction task of biomarker TP53 for TCGA-LUAD and TCGA-BLCA datasets.
@@
Ablation studies on the number of slide-level FMs. In this experiment, we conduct an ablation study on the number of slide-level FMs for distillation. We present the quantitative results of AUROC in Table 9 and Figure 8. The best prediction performance (AUROC) is obtained when 3 slide-level FMs are utilized. The performance is decreasing when the number of slide-level FMs decreases. When FuseCPath is only trained with re-embedded patch-level features, the performance decreases by 3.6% on average. These results imply that more slide-level FMs utilized for distillation will provide more useful semantic information during training.
@@ 5 Discussion and limitation
In this work, we have proposed a novel framework called FuseCPath for the fusion of multi-scale heterogeneous pathology FMs simultaneously. The proposed FuseCPath framework includes the following essential modules to effectively fuse the pathology FMs: representative patches selection based on multi-view patch features clustering, patch-level features re-embedding & aggregation, and slide-level collaborative distillation. These modules contribute to the performance improvement of biomarker prediction, gene expression, and survival analysis on datasets TCGA-LUAD, TCGA-BLCA, and TCGA-COAD. In conclusion, the FuseCPath framework will yield a new ensemble model with superior performance and benefit many meaningful downstream tasks. Additionally, clustering with multi-view features will provide insight into the visualization analysis of tissue morphography in WSI analysis.
Despite the demonstrated utility in this article, the FuseCPath poses potential limitations in its capacity to integrate more high-dimensional multi-omics data, such as the emerging spatial transcriptomic technologies. The current framework may not fully capture the underlying nonlinear relationships between different omics data. The rapid evolution of FMs presents a promising strategy for the integration of multi-omics data and WSIs [28]. In future research, we will extend the FuseCPath framework to the fusion of multi-omics representations and embeddings from multi-omics FMs to improve the precision of molecular-level WSI analysis.
@@
Zhidong Yang: Methodology, Investigation, Writing - original draft & editing. Xiuhui Shi: Conceptualization, Resources, Data curation. Wei Ba: Conceptualization, Data curation. Zhigang Song: Conceptualization, Data curation. Haijing Luan: Methodology, Validation. Taiyuan Hu: Methodology, Validation. Senlin Lin: Conceptualization, Validation. Jiguang Wang: Resources, Writing - review & editing. Shaohua Kevin Zhou: Resources, Writing - review & editing. Rui Yan: Methodology, Resources, Writing - review & editing.
Acknowledgments
This study was funded by the National Natural Science Foundation of China (62402473, 62271465), Beijing Natural Science Foundation (L252175), and the Suzhou Basic Research Program (SYG202338). J.W. lab is supported by RGC grant (R6003–22), ITC grant (ITCPD/17-9), Padma Harilela Professorship, and Sinovac Fellowship Program.